Make | Model | Problem | Solution | Date |
| | | |
Found: 776
|
Matsui | 1466 | Low, dist sound | R353 10k in audio o/p stage o/c | 14-09-04 |
Matsui | 1466 | NO PIC , SOUND , LED LIT AND CHANNEL CHANGE WORKING | R502 , R503 | 14-09-04 |
Matsui | 1466 | NO SYNC | R619 DRY | 14-09-04 |
Matsui | 1466 | PICTURE FADES TO FIELD COLL | 9v REG THERMAL OC | 14-09-04 |
Matsui | 1476 | Dead | STR50103A s/c; R2M diode s/c; 5R6 5w o/c; LOPTx also faulty | 14-09-04 |
Matsui | 1476 | Dead - no HT from STR50103A | Q502 2SA952L b-e leaky - used BC327 in lieu | 14-09-04 |
Matsui | 1476 | dead & ticking | C412 470µ 16v, C414 100µ 35v leaky & LA7837 FOP chip | 14-09-04 |
Matsui | 1476 | Field collapse | D407 s/c; R437 2R7 o/c | 14-09-04 |
Matsui | 1476 | Slow start up then blank or partially blank | R411 4k7 hi res | 14-09-04 |
Matsui | 1476 | The picture took a very long time to appear and wo | The cause of the trouble was R411 ( 4.7k ) which had risen slightly in value. | 14-09-04 |
|
Matsui | 1476 , 1492 , 2076 | REMOTES NO FUNCTIONS | CHECK IC TYPE IN TELE MAG APRIL 95 PAGE 408 | 14-09-04 |
Matsui | 1480 | Contrast control has no effect - no remote. | Standby transformer O/C. | 14-09-04 |
Matsui | 1480 | Dark pic | C509 470µ 16v | 14-09-04 |
Matsui | 1480 | DARK PICTURE | CHECK 10V LINE FOR 50HZ C509 ( 47uF ) | 14-09-04 |
Matsui | 1480 | Dead | STR50103A o/c also can be caused by R502/3 330k start up res o/c | 14-09-04 |
Matsui | 1480 | Dead | STR50103 o/c - no other faulty components | 14-09-04 |
Matsui | 1480 | Dead. | Replace STR50103 , SR2M , 2SD2105 , 6R5 7W. | 14-09-04 |
Matsui | 1480 | Int no teletext & blank raster. | Replace X901 6MHz Xtal. | 14-09-04 |
Matsui | 1480 | No raster or sound. | Replace line IC & clean the film on the nearby print. | 14-09-04 |
Matsui | 1480 | No remote or LED lights - contrast con U/S. | Standby transformer O/C. | 14-09-04 |
Matsui | 1480 | No signals | T101 pri winding o/c | 14-09-04 |
Matsui | 1480 | No sound or raster. | Replace line IC & clean the film on the nearby print. | 14-09-04 |
Matsui | 1480 | No sound or raster. | Check for sticky film around line chip. | 14-09-04 |
Matsui | 1480 | No sound/raster - no line drive | Clean sticky glue around line osc chip | 14-09-04 |
Matsui | 1480 | Partial field collapse | µ1378H field o/p chip faulty | 14-09-04 |
Matsui | 1480 | PICTURE DARK ADJUSTING A1 CONTROL HELPS BUT STILL | C509 470uF 16v | 14-09-04 |
Matsui | 1480 | STILL DEAD AFTER REPLACING D508 ( R2M ) , STR50103 A | Q108 | 14-09-04 |
Matsui | 1480 | STUCK IN STANDBY | C107 DRY | 14-09-04 |
Matsui | 1480A | Dark pic with no brightness control | C483 4µ7 50v in beam lim | 14-09-04 |
Matsui | 1480A | Dead after replacement of usual components | Q108 2SB698E power sw tr connected to pin 2 IC501 leaky. Can use BC327 in lieu | 14-09-04 |